A grocery store sells 3 pounds of bananas and 2 pounds of apples for $2.40. The cost of 4 pounds of bananas and 5 pounds of apples is $ 4.60. Find the cost of 1p ound of apples and 1 pound of bananas
Answer by mananth(16946)

x= cost of one pound bananas
y= cost of one pound apples
Total sold
3.00 x + 2.00 y = 2.40 .............1
Total value
4.00 x + 5.00 y = 4.60 .............2
Eliminate y
multiply (1)by -5.00
Multiply (2) by 2.00
-15.00 x -10.00 y = -12.00
8.00 x + 10.00 y = 9.20
Add the two equations
-7.00 x = -2.80
/ -7.00
x = 0.40
plug value of x in (1)
3.00 x + 2.00 y = 2.40
1.20 + 2.00 y = 2.40
2.00 y = 2.40 -1.20
2.00 y = 1.20
y = 0.60
x= 0.40 cost of one pound bananas
y= 0.60 cost of one pound apples
m.ananth@hotmail.ca